The density of a hydrocarbon gas (alkynes group) reported to be 1.7847 g/L at STP, identify (i.e. give the name and chemical formula) this gas?
1
Expert's answer
2012-11-30T07:11:20-0500
& 1.7847 1
CxHy ------ & CxHy
& Mw & 22.4

Mw=1.7847*22.4=40

40/12=3.333& so x=3

40-12*3= 8 y=8

Answer& C3H8